# SFY — A lightweight, high-frequency and phase-resolving wave-buoy for coastal waters

### v2

@@ 1 Introduction
Of the many ways to measure waves, in situ heave buoys and wave staffs are the most popular (Holthuijsen, 2010). There exists a range of high–quality commercial moored and drifting wave buoys [see, e.g., the buoys by Datawell (Datawell BV, 2023), SOFAR Ocean (Raghukumar et al., 2019), and Aanderaa (Saetre et al., 2023), LainePoiss (Alari et al., 2022)]. These buoys are designed for operational application and therefore have a need to function robustly for long stretches of times. They usually report statistical properties, like significant wave height, and in some cases time series up to a few Hz. Often the time series are only available onboard, and not over telemetry. However, the last decade has seen a small revolution in low-cost, open-source wave-measuring devices (Thomson et al., 2023; Rabault et al., 2020; Feddersen et al., 2023a; Collins et al., 2024). These designs generally do not seek to have the same robustness as commercial buoys, but can measure in quite different conditions. Because of the low cost and lower environmental impact, they are also often considered expendable. This is important when measuring in regions where the risk of loss and damage is high. More importantly, these instruments are typically one or more orders of magnitude smaller and lighter than traditional wave buoys and can thus measure in locations where traditional buoys cannot. Crucially, they also offer the opportunity to measure quantities that larger buoys are incapable of capturing, such as the acceleration in breaking waves (Feddersen et al., 2023b). The limits and capabilities of the instrument determine which questions can be answered. Heterogeneous and different solutions are valuable in themselves in that they have different biases, cover different use cases and make results more robust.
A relevant example to this work is Brown et al. (2018) who previously quantified breakers in the open ocean by measuring accelerations. Similarly, the microSWIFT (Thomson, 2012; Thomson et al., 2019) has been used to study waves near the shore and in areas with strong currents (Rainville, 2022; Rainville et al., 2023). Sinclair (2014) and later Brown and Paasch (2021) and Fisher et al. (2021); Feddersen et al. (2023a) have attempted to measure breaking waves in the surf or in wave laboratories.
@@
The main improvements of the SFY buoy is that it allows us to transmit much more data, is lightweight, and low cost with a small physical and environmental footprint. This combination provides a number of advantages: It can transmit continuous high-frequency time series of the full three-dimensional acceleration vector measured by the buoy. It can be deployed easily from surf-boards or small vessels, or in lightweight moorings. It can be deployed in arrays of many buoys capturing the spatial variability, which is perhaps more important than a single high–quality point measurement in coastal environments. And it can resolve the phase, impulses and trajectory of waves and breaking waves, also between buoys in the same dense spatial network. The buoy does not currently have a magnetometer, but tracks the orientation in its own reference system on short time scales. The deployments must be within the cellular coverage and will not work far from the coast unless there is a marine network available.
@@
Here, $\omega=2\pi f$ is the angular frequency [rad/s], $f$ is the linear frequency [Hz] and $\mathcal{F}$ is the Fourier transform. When analyzing spectra of elevation energy (variance density spectrum), $E$, no detrending and filtering is necessary. The integration of the acceleration density ($A$), Eq. (2), is performed in one step:
@@
The SFY buoy comes in a small free-drifting form and in a larger moored form. The free-drifting buoy has a heave response of approximately $f_{\mathrm{hr}}=3.56$ Hz, while the moored buoy has $f_{\mathrm{hr}}=1.44$ Hz (assuming the buoy is 50 % submerged). Clearly, though, the expression for the heave response is an approximation, and does not take the horizontal length of the buoy into account. At a rule-of-thumb level, one can argue that if the wavelength ($\lambda$) of the wave to be measured is shorter than twice the (horizontal) radius of the buoy, $r$, the buoy will not be able to follow the wave. In deep water, the linear phase speed is $c_{\mathrm{p}}=\sqrt{g/k}$ with $k=2\pi/\lambda$ the wavenumber. From this we find that the geometric cut-off frequency ($f_{\mathrm{gc}}$) where the wavelength equals $2r$ is
@@ 2 Buoy design
Figure 1: The components that make up the SFY wave buoy. From the left: housing (plastic bottle) with foam for additional stability, battery pack (alkaline), main board shown twice (both sides): Microcontroller Unit (MCU) in red, modem and GPS in green, SD-card, IMU, and GPS and cellular antennas. For scale: the ruler is 60 cm long.
@@ 3 Lab experiment: Wave flume
Experiments were carried out in a wave flume at SINTEF Ocean in Trondheim, Norway (described in Singsaas et al. (2020)). The flume is 12 m long and 0.53 m wide, with a piston-type wave maker from Edinburgh Designs (Edinburgh Designs, 2023) at one end, and a perforated absorbing beach at the other end. The wave maker is programmable and can be set to generate simple harmonic waves as well as wave packets and random realisations of a specified wave spectrum.
@@ 3.1 Significant wave height
1.00 1 0.05 / 0.14 0.15 9.60 0.15
1.50 7a 0.03 / 0.08 0.08 4.68 7.11 4.23 2.14
1.50 7b 0.03 / 0.08 0.08 3.43
1.50 8a 0.02 / 0.06 0.08 1.78
Table 2: Significant wave height measured in the wave flume, automatic high-pass cut-off frequency.
@@
[…]
$H_{m0}$ $=4\cdot\sqrt{\frac{1}{2}A_{p}^{2}}$ (7b)
The measured signal from the buoys need to be high-pass filtered at a cut-off frequency to avoid low-frequency noise. The wave gauge is not filtered, since it has a flat noise curve on displacement (Fig. 2). Automatic selection of low-end cut-off frequency is generally hard. This is a problem that affects IMU based buoys in general (see, e.g., Nose et al. (2018)) and will be discussed further in Section 66.1. Alari et al. (2022) chose the bandpass cut-off frequencies of the measured signal in the wave tank tests to be $0.30$ and $1.28$ Hz and the LainePoiss® buoy was evaluated at frequencies between $0.57$ and $0.73$ Hz.
In Table 2 the cut-off frequency is found automatically by searching for the first spectral minimum above $0.05$ Hz. The algorithm is based on the process outlined in Rabault et al. (2022) (Figure 7) as well as in Tucker (1958); Waseda et al. (2018). However, when this algorithm is unsuccessful it results in unphysically large amplitudes and significant wave height estimates. The auto-detected cut-off frequency is shown and listed in Fig. 2. For instance for $1$ Hz in Table 2 and Fig. 2f the significant wave height for B23 is obviously wrong and overestimated compared to WOB1. This can be attributed to the cut-off frequency being chosen too low for B23, while better for WOB1 (see $f_{c}$ in Fig. 2f, note that the cut-off for B23 is off the scale of the plot). In general it is easier to auto-detect the cut-off frequency when the waves are higher and the signal to noise ratio is greater. Since we know the input frequency for these experiments we can manually choose cut-off frequencies. In Table 3 the significant wave height is shown for the manually chosen frequency band. These values match better, and do not show exaggerated significant wave heights like those present in Table 2 (where a too low or too high cut-off frequency has been picked). In unknown conditions it is therefore best to start with a minimum of $0.05$ Hz for this buoy, and possibly other IMU based buoys. If the wave height suddenly jumps to unrealistic levels or there is significant energy at very low frequencies this is a sign that it should be increased, or that the measurements have been compromised in other ways (e.g. a poor mooring, unstable float or interference from traffic).
@@ 3.2 Wave buoy orbital measurements
In order to obtain the trajectory of the wave buoy we apply principal component analysis (PCA) to the high-pass filtered $x$ and $y$ acceleration. The acceleration is high-pass filtered (using the Butterworth-filter) so that only movement due to wave motion. The first principal component, $\bf u_{0}$, is the direction in the $xy$-plane along which the variation is greatest, the second component ($\bf u_{1}$) is orthogonal to $\bf u_{0}$ representing the rest of the variance.
@@
[|$\ \bf h_{i}\\ |\end{bmatrix}=\begin{bmatrix}|&|\\ \bf x&\bf y\\ |&|\end{bmatrix}\cdot\frac{\bf u_{i}}{\left\|\bf u_{i}\right\|}$ (8)
$\bf h_{0}$ and $\bf h_{1}$ are the vectors with re-projected $x$ and $y$ accelerations.
If the buoy is only moving along one direction, and the gyroscope is not drifting, all movement should be aligned along the first principal component. It is clearly visible in Fig. 4 that this is the case for almost all the movement. The ratio between the two principal components for 10 seconds in the wave-flume is $925$ (upper panel in Fig. 4), which provides a measure of how robust this method is on this time scale.
@@
Linear waves predict the orbital motion of particles. If the wave-buoy is to measure the details of wave-motion it should be water-following. The relation between the horizontal and vertical motion, and their spectra, will tell us how well the buoy is following the particle motion in a wave (Feddersen et al., 2023a). The amplitudes of the horizontal and vertical particle orbital velocity in a plane wave are (e.g. Holthuijsen (2010))
@@ 4 Open water: Fedjeosen
Figure 8: Observations from the buoy compared to the simulated $H_{m0}$ (WAM800 with current). Linear regression and quantile-quantile (Q-Q) lines are shown. The model predict slightly higher significant wave height, but with a otherwise fairly linear relationship (Pearson correlation coefficient, $R=0.96$). The model without current shows the same $R$-value.
@@
The measured significant wave height is compared to the output of significant wave height from the WAM domain for the west coast of Norway, Vestlandet (Fig. 7 and 8), with and without currents. The nearest grid point in the wave model is about $280$ m away from the buoy location. Figure 8 compares the observations against the simulated $H_{m0}$, and there is a good match with an $R$-value of $0.96$ for both models. The buoy measures slightly lower $H_{m0}$ values than model predicts, except for a few instances of higher wave heights where the buoy measures higher than the model predicts. In general the relationship is close to linear.
@@
Figure 9: Elevation energy (variance density) spectra for comparable situations (similar $H_{m0}$ and $T_{p}$) of the same buoy (Fig. 6) deployed at two different locations. The IMU (before feeding the AHRS-algorithm) was configured to sample at $52$ Hz in the first case (Hjeltefjorden), and at $208$ Hz in the second case (Fedjeosen). The arrow in the lower left panel shows low-frequency noise caused by insufficient sample rate.
@@
$A(f,t^{\prime})=|\mathcal{F}\{a_{z}\}|+|\mathcal{F}\{a_{x}\}|+|\mathcal{F}\{a_{y}\}|\quad\textrm{for time-windows}\;\left<t^{\prime},t^{\prime}+t_{bw}\right>$ (12)
@@
A way to quantify breaking waves is to estimate the breaking frequency (or approximately breaking probability) (Schwendeman and Thomson, 2014; Brown et al., 2018),
@@
Here, $N_{b}$ is the number of breakers detected during a measurement interval, $\tau=20$ min, and $T$ is the wave period, here approximated with $T_{m01}$. Figure 10a shows $Q_{b}$ (lower panel) for a subset of the data. We have visually inspected images taken every 15 minutes with an air traffic camera on the Fedje Island pointing in the direction of the wave buoy. In Fig. 10 photos from three instants are shown. These are marked in Fig. 10a with vertical dashed lines. Qualitatively, the detection of breakers agrees well with the visual impression of breaking waves in the photographs. The wave steepness ($H_{m0}/(g\cdot\frac{T_{m01}^{2}}{2\pi})$) for the measurement period varies very little (but agrees well with models).
@@ 5.2 Trajectory in a breaking wave
Figure 14: Spectrum captured outside the surf at Green Bowl, Bali.
@@
The recorded displacement amplitudes match well with expected values from the spectrum and visual observations. Usually, measurements from breaking waves in the surf are not suitable to be integrated to displacement without heavy filtering. In Fig. 16 we can match the photographs with the integrated trace (Fig. 15), and see that the recorded trajectory (Fig. 16d) is sensible even with a very low cut-off frequency. This is discussed further in Sec. 66.2.
@@ 6.1 Low-frequency noise and breaking waves
All IMU-based wave buoys will struggle to measure low-frequency signals. The sensor measures acceleration, and the double integration to obtain vertical displacement ((1a) and (1b)) introduces drift which must be removed using detrending. Additionally, a core assumption of AHRS-algorithms is that they measure a high-frequency signal, and very long period waves may not fall into this domain and appear to shift the earth’s gravity vector from the buoys point of view.
@@
Most other assessments of the performance and behavior of IMU based buoys observe and identify the low-frequency noise (see references above), but its origin is not well explained. Some of it is attributed to the integrated noise of the IMU. However, that does not explain why there is much higher low-frequency noise in some situations than others, or that the noise is well above the expected noise-floor (when extrapolating the noise curve from high to low-frequencies). Alari et al. concluded that the low-frequency energy is misplaced from other frequencies, and that it should not be filtered completely away for calculation of spectral moments as is the usual way to deal with this noise. Brown et al. (2018) also attribute some of the low-frequency noise as alias from higher frequencies. For the MOTUS buoy (Dorgeville et al., 2018), a mechanical filter was used to reduce measured vibrations (from wind, waves, and currents) above the Nyquist-frequency that fold back into lower frequencies. However, for the SFY both the gyroscope and accelerometer have built-in low-pass filters that would avoid sampling alias.
@@
We observe that the low-frequency noise occurs more frequently in the surf, and only sporadically in the open ocean. As wave breaking is more violent and happens much more often in the surf zone, it seems likely that high-frequency acceleration impulses from breaking waves are the cause of the low-frequency noise in the measured elevation. Specifically, rapid acceleration and rotation caused by a breaking wave insufficiently sampled into the fusion algorithm would cause the buoy to lose orientation and would no longer correctly extract the vertical acceleration. To test this hypothesis we configured buoys with different IMU sample rates ($208$ Hz and $52$ Hz), but equal output rate (filtered using a FIR-filter and decimated). In the wave flume (Sec. 3), both the low-frequency and the high-frequency buoys performed well on harmonic waves (Fig. 3). However, when the buoys are subjected to breaking waves in the wave flume (Sec. 33.3), the buoy with the higher IMU-sampling (B25) performs far better than the other (B23), with low-frequency oscillations clearly visible.
@@
A buoy that encounters breaking waves only sporadically yield a spectrum contaminated by these impulses. If this is the case, higher sample rate should better resolve the rotation and acceleration and thus result in less low-frequency noise. The spectra in Fig. 9 suggest that the buoy with $208$ Hz sampling and AHRS frequency performs better on low-frequencies than the one with $52$ Hz sample rate. Detecting and removing segments with breaking impulses from the time series should also improve the spectral estimates.
@@
In Sec. 5, Fig. 16d we were able to retrieve the elevation with a cut-off frequency as low as $0.04$ Hz.

### v3

@@ 1 Introduction
Of the many ways to measure waves, in situ heave buoys and wave staffs are the most popular (Holthuijsen, 2010). There exists a range of high-quality commercial moored and drifting wave buoys [see, e.g., the buoys by Datawell (Datawell BV, 2023), SOFAR Ocean (Raghukumar et al., 2019), Aanderaa (Saetre et al., 2023), and LainePoiss (Alari et al., 2022)]. These buoys are designed for operational application and therefore have a need to function robustly for long stretches of times. They usually report statistical properties, like significant wave height, and in some cases time series up to a few Hz. Often the time series are only available onboard, and not over telemetry. However, the last decade has seen a small revolution in low-cost, open-source wave-measuring devices (Thomson et al., 2023; Rabault et al., 2020; Feddersen et al., 2023a; Collins et al., 2024). These designs generally do not seek to have the same robustness as commercial buoys, but can measure in quite different conditions. Because of the low cost and lower environmental impact, they are also often considered expendable. This is important when measuring in regions where the risk of loss and damage is high. More importantly, these instruments are typically one or more orders of magnitude smaller and lighter than traditional wave buoys and can thus measure in locations where traditional buoys cannot. Crucially, they also offer the opportunity to measure quantities that larger buoys are incapable of capturing, such as the acceleration in breaking waves (Feddersen et al., 2023b). The limits and capabilities of the instrument determine which questions can be answered. Heterogeneous and different solutions are valuable in themselves in that they have different biases, cover different use cases and make results more robust.
A relevant example to this work is Brown et al. (2018) who previously quantified breakers in the open ocean by measuring accelerations. Similarly, the microSWIFT (Thomson, 2012; Thomson et al., 2019) has been used to study waves near the shore and in areas with strong currents (Rainville, 2022; Rainville et al., 2023). Sinclair (2014) and later Brown and Paasch (2021), Fisher et al. (2021), and Feddersen et al. (2023a) have attempted to measure breaking waves in the surf or in wave laboratories.
@@
The main improvements of the SFY buoy is that it allows us to transmit much more data, is lightweight, and low cost with a small physical and environmental footprint. This combination provides a number of advantages: It can transmit continuous high-frequency time series of the full three-dimensional acceleration vector measured by the buoy. It can be deployed easily from surf-boards or small vessels, or in lightweight moorings. It can be deployed in arrays of many buoys capturing the spatial variability, which is perhaps more important than a single high-quality point measurement in coastal environments. And it can resolve the phase, impulses and trajectory of waves and breaking waves, also between buoys in the same dense spatial network. The buoy does not currently have a magnetometer, but tracks the orientation in its own reference system on short time scales. The deployments must be within the cellular coverage and will not work far from the coast unless there is a marine network available.
@@
Here, $\omega=2\pi f$ is the angular frequency [rad/s], $f$ is the linear frequency [Hz] and $\mathcal{F}$ is the Fourier transform. When analyzing spectra of elevation energy (variance density spectrum), $E$, no detrending and filtering is necessary. The integration of the acceleration density ($A$, found using Welch’s method), Eq. (2), is performed in one step:
@@
The SFY buoy has been deployed in a small free-drifting form and in a larger moored form. The free-drifting buoy has a heave response of approximately $f_{\mathrm{hr}}=3.56$ Hz, while the moored buoy has $f_{\mathrm{hr}}=1.44$ Hz (assuming the buoy is 50 % submerged). Clearly, though, the expression for the heave response is an approximation, and does not take the horizontal length of the buoy into account. At a rule-of-thumb level, one can argue that if the wavelength ($\lambda$) of the wave to be measured is shorter than twice the (horizontal) radius of the buoy, $r$, the buoy will not be able to follow the wave. In deep water, the linear phase speed is $c_{\mathrm{p}}=\sqrt{g/k}$ with $k=2\pi/\lambda$ the wavenumber. From this we find that the geometric cut-off frequency ($f_{\mathrm{gc}}$) where the wavelength equals $2r$ is
@@ 2 Buoy design
Figure 1: The components that make up the SFY wave buoy. From the left: housing (plastic bottle) with foam for additional stability, battery pack (alkaline), main board shown twice (both sides): Microcontroller Unit (MCU) in red, modem and GPS in green, SD-card, IMU, and GPS and cellular antennas. For scale: the ruler is 85 cm long.
@@ 3 Lab experiment: Wave flume
Experiments were carried out in a wave flume at SINTEF Ocean in Trondheim, Norway (described in Singsaas et al. (2020)). The flume is 12 m long and 0.53 m wide, with a piston-type wave maker from Edinburgh Designs (Edinburgh Designs, 2023) at one end, and a perforated absorbing beach at the other end. The wave maker is programmable and can be set to generate simple harmonic waves as well as wave packets and random realisations of a specified wave spectrum. In the experiment below it is configured to generate harmonic waves of different amplitudes and frequencies, as well as focused breaking waves.
@@ 3.1 Significant wave height
1.00 1 0.05 / 0.14 0.15 $9.60^{\dagger}$ 0.15
1.50 7a 0.03 / 0.08 0.08 $4.68^{\dagger}$ $7.11^{\dagger}$ $4.23^{\dagger}$ $2.14^{\dagger}$
1.50 7b 0.03 / 0.08 0.08 $3.43^{\dagger}$
1.50 8a 0.02 / 0.06 0.08 $1.78^{\dagger}$
Table 2: Significant wave height measured in the wave flume, automatic high-pass cut-off frequency. Values marked with $\dagger$ are unreasonably large because the automatic cut-off frequency is chosen to low, and includes low-frequency noise.
@@
$H_{m0}=4\sqrt{m_{0}}=4\sqrt{\frac{1}{2}A_{p}^{2}}$ (7)
The measured signal from the buoys need to be high-pass filtered at a cut-off frequency, $f_{c}$, to avoid low-frequency noise. The wave gauge is not filtered, since it has a flat noise curve on displacement (Fig. 2). Automatic selection of low-end cut-off frequency is generally hard. This is a problem that affects IMU based buoys in general (see, e.g., Nose et al. (2018)) and will be discussed further in Section 66.1. Alari et al. (2022) chose the bandpass cut-off frequencies of the measured signal in the wave tank tests to be $0.30$ and $1.28$ Hz and the LainePoiss® buoy was evaluated at frequencies between $0.57$ and $0.73$ Hz.
In Table 2 the cut-off frequency is found automatically by searching for the first spectral minimum above $0.05$ Hz using the find_peaks algorithm in SciPy [^3] https://github.com/gauteh/sfy/blob/v1.0/sfy-processing/sfy/signal.py#L319 (minimum distance of $7$ frequency bins for $52$ Hz sample rate, $3$ for $26$ Hz sample rate, and a prominence of $0.05$). The algorithm is based on the process outlined in Rabault et al. (2022) (Figure 7) as well as in Tucker (1958); Waseda et al. (2018). However, when this algorithm is unsuccessful it results in unphysically large amplitudes and significant wave height estimates. The auto-detected cut-off frequency is shown and listed in Fig. 2. For instance for $1$ Hz in Table 2 and Fig. 2f the significant wave height for B23 is obviously wrong and overestimated compared to WOB1. This can be attributed to the cut-off frequency being chosen too low for B23, while better for WOB1 (see $f_{c}$ in Fig. 2f, note that the cut-off for B23 is off the scale of the plot). In general it is easier to auto-detect the cut-off frequency when the waves are higher and the signal to noise ratio is greater. Since we know the input frequency for these experiments we can manually choose cut-off frequencies. In Table 3 the significant wave height is shown for the manually chosen frequency band. These values match better, and do not show exaggerated significant wave heights like those present in Table 2 (where a too low or too high cut-off frequency has been picked). In unknown conditions it is therefore best to start with a minimum of $0.05$ Hz for this buoy, and possibly other IMU based buoys. If the wave height suddenly jumps to unrealistic levels or there is significant energy at very low frequencies this is a sign that it should be increased, or that the measurements have been compromised in other ways (e.g. a poor mooring, unstable float or interference from traffic).
@@ 3.2 Wave buoy orbital measurements
In order to obtain the trajectory of the wave buoy we apply principal component analysis (PCA) to the high-pass filtered $x$ and $y$ acceleration. The acceleration is filtered (using the Butterworth-filter) so that only movement due to wave motion is included. The first principal component, $\bf u_{0}$, is the direction in the $xy$-plane along which the variation is greatest, the second component ($\bf u_{1}$) is orthogonal to $\bf u_{0}$ representing the rest of the variance.
@@
$\begin{bmatrix}h_{i}^{1}\\ \vdots\\ h_{i}^{N}\end{bmatrix}=\begin{bmatrix}x^{1}&y^{1}\\ \vdots&\vdots\\ x^{N}&y^{N}\end{bmatrix}\cdot\frac{\bf u_{i}}{\left\|\bf u_{i}\right\|}$ (8)
$h_{0}^{n}$ and $h_{1}^{n}$ are the vector components with re-projected $x$ and $y$ accelerations, and the superscript index $n$ indicates time from $t_{1}$ to $t_{N}$.
If the buoy is only moving along one direction, and the gyroscope is not drifting, all movement should be aligned along the first principal component. It is clearly visible in Fig. 4 that this is the case for almost all the movement. The ratio between the two principal components (variances) for 10 seconds in the wave-flume is $925$ (upper panel in Fig. 4), which provides a measure of how robust this method is on this time scale.
@@
Linear wave theory predicts the orbital motion of particles. If the wave-buoy is to measure the details of wave-motion it should be water-following. The relation between the horizontal and vertical motion, and their spectra, will tell us how well the buoy is following the particle motion in a wave (Feddersen et al., 2023a). The amplitudes of the horizontal and vertical particle orbital velocity in a plane wave are (e.g. Holthuijsen (2010))
@@ 4 Open water: Fedjeosen
Figure 8: Observations from the buoy compared to the simulated $H_{m0}$ (WAM800 with current). Linear regression and quantile-quantile (Q-Q) lines are shown. The model predicts slightly higher significant wave height, but with an otherwise fairly linear relationship (Pearson correlation coefficient, $R=0.96$). The model without current shows the same $R$-value.
@@
The measured significant wave height is compared to the output of significant wave height from the WAM domain for the west coast of Norway, Vestlandet (Fig. 7 and 8), with and without currents. The nearest grid point in the wave model is about $280$ m away from the buoy location. Figure 8 compares the observations against the simulated $H_{m0}$, and there is a good match with an $R$-value of $0.96$ for both models. The buoy measures slightly lower $H_{m0}$ values than the model predicts, except for a few instances of higher wave heights where the buoy measures higher than the model predicts. In general the relationship is close to linear.
@@
Figure 9: Elevation energy (variance density) spectra for comparable situations (similar $H_{m0}$ and $T_{p}$) of the same buoy (Fig. 6) deployed at two different locations. The IMU (before feeding the AHRS-algorithm) was configured to sample at $52$ Hz in the first case (Hjeltefjorden), and at $208$ Hz in the second case (Fedjeosen). The arrow in the lower left panel shows low-frequency noise common to IMU buoys caused by insufficient sampling of impulsive events (see Sec. 66.1 for an extensive discussion).
@@
$A(f,t^{\prime})=|\mathcal{F}\{a_{z}\}|+|\mathcal{F}\{a_{x}\}|+|\mathcal{F}\{a_{y}\}|$ (12)
@@
A way to quantify breaking waves is to estimate the breaking fraction, the fraction of waves passing a point that are breaking (or approximately breaking probability) (Schwendeman and Thomson, 2014; Brown et al., 2018),
@@
Here, $N_{b}$ is the number of breakers detected during a measurement interval, $\tau=20$ min, and $T$ is the wave period, here approximated with $T_{m01}$. Figure 10a shows $Q_{b}$ (lower panel) for a subset of the data. We have visually inspected images taken every 15 minutes with an air traffic camera on the Fedje Island pointing in the direction of the wave buoy. In Fig. 10 photos from three instants are shown. These times are marked in Fig. 10a with vertical dashed lines. Qualitatively, the detection of breakers agrees well with the visual impression of breaking waves in the photographs. The wave steepness ($H_{m0}/(g\cdot\frac{T_{m01}^{2}}{2\pi})$) for the measurement period varies very little (but agrees well with models).
@@ 5.2 Trajectory in a breaking wave
Figure 14: Spectrum captured outside the surf at Green Bowl, Bali. The vertical line shows the high-pass cut-off frequency ($f_{c}$) to avoid the low-frequency noise common to IMU buoys (see Sec. 66.1 for an extensive discussion).
@@
The recorded displacement amplitudes match well with expected values from the spectrum and visual observations. Usually, measurements from breaking waves in the surf are not suitable to be integrated to displacement without heavy filtering. In Fig. 16 we can match the photographs with the integrated trace (Fig. 15), and see that the recorded trajectory (Fig. 16d) is sensible even with a very low cut-off frequency. We are at the limit of what trajectories are possible to reconstruct, therefore the trajectory may be tilted due to insufficient speed of the AHRS-algorithm. The time of the images may also be slightly shifted, as synchronization is insufficiently accurate. This is discussed further in Sec. 66.2.
@@ 6.1 Low-frequency noise and breaking waves
All IMU-based wave buoys will struggle to measure low-frequency signals. The sensor measures acceleration, and the double integration to obtain vertical displacement (Eqs. (1a) and (1b)) introduces drift which must be removed using detrending. Additionally, a core assumption of AHRS-algorithms is that they measure a high-frequency signal, and very long period waves may not fall into this domain and appear to shift the Earth’s gravity vector from the buoys point of view.
@@
Most other assessments of the performance and behavior of IMU based buoys observe and identify the low-frequency noise (see references above), but its origin is not well explained. Some of it is attributed to the integrated noise of the IMU. However, that does not explain why there is much higher low-frequency noise in some situations than others, or that the noise is well above the expected noise-floor (when extrapolating the noise curve from high to low-frequencies). Alari et al. concluded that the low-frequency energy is misplaced from other frequencies, and that it should not be filtered completely away for calculation of spectral moments as is the usual way to deal with this noise. Brown et al. (2018) also attribute some of the low-frequency noise as aliasing from higher frequencies. For the MOTUS buoy (Dorgeville et al., 2018), a mechanical filter was used to reduce measured vibrations (from wind, waves, and currents) above the Nyquist-frequency that fold back into lower frequencies. However, for the SFY both the gyroscope and accelerometer have built-in low-pass filters that would avoid sampling alias.
@@
We observe that the low-frequency noise occurs more frequently in the surf, and only sporadically in the open ocean. As wave breaking is more violent and happens much more often in the surf zone, it seems likely that high-frequency acceleration impulses from breaking waves are the cause of the low-frequency noise in the measured elevation. Specifically, rapid acceleration and rotation caused by a breaking wave insufficiently sampled into the fusion algorithm would cause the buoy to lose orientation and no longer correctly extract the vertical acceleration. To test this hypothesis we configured buoys with different IMU sample rates ($208$ Hz and $52$ Hz), but equal output rate (filtered using a FIR-filter and decimated). In the wave flume (Sec. 3), both the low-frequency and the high-frequency buoys performed well on harmonic waves (Fig. 3). However, when the buoys are subjected to breaking waves in the wave flume (Sec. 33.3), the buoy with the higher IMU-sampling (B25) performs far better than the other (B23), with low-frequency oscillations clearly visible.
@@
A buoy that encounters breaking waves only sporadically yields a spectrum contaminated by these impulses. If this is the case, higher sample rate should better resolve the rotation and acceleration and thus result in less low-frequency noise. The spectra in Fig. 9 suggest that the buoy with $208$ Hz sampling and AHRS frequency performs better on low-frequencies than the one with $52$ Hz sample rate. Detecting and removing segments with breaking impulses from the time series should also improve the spectral estimates.